# MealRec: A Meal Recommendation Dataset with Meal-Course Affiliation for Personalization and Healthiness

### v1

@@ Abstract.
Meal recommendation, as a typical health-related recommendation task, contains complex relationships between users, courses, and meals. Among them, meal-course affiliation associates user-meal and user-course interactions. However, an extensive literature review demonstrates that there is a lack of publicly available meal recommendation datasets including meal-course affiliation. Meal recommendation research has been constrained in exploring the impact of cooperation between two levels of interaction on personalization and healthiness. To pave the way for meal recommendation research, we introduce a new benchmark dataset called MealRec. Due to constraints related to user health privacy and meal scenario characteristics, the collection of data that includes both meal-course affiliation and two levels of interactions is impeded. Therefore, a simulation method is adopted to derive meal-course affiliation and user-meal interaction from the user’s dining sessions simulated based on user-course interaction data. Then, two well-known nutritional standards are used to calculate the healthiness scores of meals. Moreover, we experiment with several baseline models, including separate and cooperative interaction learning methods. Our experiment demonstrates that cooperating the two levels of interaction in appropriate ways is beneficial for meal recommendations. Furthermore, in response to the less healthy recommendation phenomenon found in the experiment, we explore methods to enhance the healthiness of meal recommendations. The dataset is available on GitHub (https://github.com/WUT-IDEA/MealRecPlus).
@@ 1. Introduction
To pave the way for meal recommendation studies, we introduce a meal recommendation benchmark dataset called MealRec. Specifically, considering the category-constraints feature of meals, a mapping from course name phrases to course categories is established. Then, inspired by research on conversational recommender systems, meal-course affiliation and user-meal interaction are derived from user’s dining sessions simulated based on user-course interaction data from a recipe-sharing website222https://www.allrecipes.com. Based on the principle of collaborative filtering that similar users may have similar behaviors, user-meal interactions are mined to expand the scale of data. Finally, two well-known nutritional standards from the World Health Organization and the United Kingdom Food Standards Agency are used to evaluate healthiness of meals. The healthiness analysis of the dataset finds that 1) users have different preferences for healthiness at course-level and meal-level, and 2) users prefer less healthy meals. To explore the impact of different separate and cooperative interaction learning methods on meal recommendation, we experiment with several state-of-the-art models and establish baseline performances. Our experiment demonstrates that cooperating the two levels of interaction in appropriate ways is beneficial for meal recommendations. Furthermore, in response to the less healthy recommendation phenomenon found in the experiment, we explore potential methods to improve the healthiness of recommendation results. In the end, we point out possible directions for building personalized and healthy meal recommender systems, with the long-term goal of transforming how users choose food conveniently and healthily. This aligns with AI for good, viz., utilizing technology’s potential to promote public good health and well-being (Floridi et al., 2021a; Tomašev et al., 2020; Floridi et al., 2021b).
Our main contributions lie two-fold: 1) We introduce a new meal recommendation benchmark dataset, MealRec, which fills a data gap by including meal-course affiliation and healthiness information. It provides conditions for researching cooperative interaction learning methods on meal recommendation, and serves as a valuable resource for AI for Good research in the food domain. 2) Our experimental study investigates separate and cooperative interaction learning methods and potential ways for improving the healthiness in meal recommendation, which provides useful insights into diet-based public health promotion.
@@ 2.2. Meal Recommendation Datasets
In general, an extensive literature review demonstrates a lack of publicly available meal recommendation datasets with meal-course affiliation. The main reason is that the two main data collecting methods are hindered due to user health privacy and meal scenario characteristics. To create conditions for advancing representation learning in meal recommendation research, we construct a new meal recommendation dataset called MealRec. A simulation approach is adopted to construct MealRec, which is inspired by the method of simulating user dialogues to build datasets in conversational recommender system research (Sun and Zhang, 2018; Qu et al., 2023; Tu et al., 2022).
@@ 3. Construction of the Dataset
In order to find the base dataset for constructing our dataset, we search extensively in the literature and the Internet. Finally, we find two open-source large-scale food recommendation datasets, foodRecSys-V1444https://www.kaggle.com/elisaxxyga/foodrecsysv1 and Food180K555https://www.kaggle.com/datasets/shuyangli94/food-com-recipes-and-user-interactions, which come from two very popular food sharing websites666https://www.food.com respectively. These two datasets have similar characteristics, such as having similar data sizes and rich course description information. The main difference is that the data density of foodRecSys-V1 is three times that of Food180K, and Food180K contains category tags, but foodRecSys-V1 does not. In our practice, data density is crucial to the construction of meal datasets, so we finally choose foodRecSys-V1 as our base dataset and use the course category tags in Food180K to build a mapping from course name to course category. foodRecSys-V1 contains 49,698 courses and 3,794,003 user-course interactions. The structure of the base dataset is shown in Table 2. The course data contains rich descriptive information including nutrition. The interaction data includes timestamps of each user’s reviews on courses.
@@
| timestamp | the time last modified | 2014-04-25T14:54:20 | ||
@@ 3.1. Course Category Mapping
Based on the mapping table between course name phrases and categories, some courses in the foodRecSys-V1 dataset are assigned categories. Before constructing meal data, we need to clarify the composition of a meal. The basic meal comprises three courses: an appetizer, a main dish, and a dessert. In this paper, we refer to the composition of the three-course meal as these three categories have the most courses in the base dataset.
@@ 3.4. Healthiness Measuring
The courses in the base datatset come from the Allrecipes.com website and contain detailed nutritional data powered by the ESHA Research Database 777https://esha.com/. The data format of nutritional data is shown in Figure 3. These nutritional data, quantified per serving, cover all macro-nutrients mentioned in two nutritional standards above. We calculate both the FSA score and WHO score for each course. Following other food-related studies (Trattner and Elsweiler, 2017b), the healthiness of a meal is defined as the mean of FSA/WHO scores of its courses. Thus far, we have constructed the dataset, which called as MealRec.
@@ 3.5.1. Data Statistics
Compared with other meal recommendation datasets in Table 1, MealRec has both levels of interaction and meal-course affiliation, which can support the research of cooperative interaction learning methods. A bundle is a set of items as a whole (Zhu et al., 2014), such as a clothing outfit, a book list, a music playlist, etc. A meal is a type of bundle in the food domain. Research on bundle recommendation (Chang et al., 2020; Ma et al., 2022; Ke et al., 2023a) has emerged in recent years, and Clothing (Sun et al., 2022) and Youshu (Chen et al., 2019) are important benchmark datasets. However, they can’t be applied directly to meal recommendations due to the lack of healthiness information. Table 4 shows the comparison in data size between them. Since Youshu is one of the most important datasets of the same type, we intentionally construct the data size of MealRec to be comparable to Youshu. Specifically, the length of the simulated user session affects the size of the constructed dataset, which is chosen to start with a short time and gradually extend until the data size is comparable to Youshu. Furthermore, Clothing and Youshu are used by existing bundle recommendation research to experiment with data-driven deep learning models. Therefore, this comparison of data size shows that MealRec can support the applications of existing advanced recommendation technologies to meal recommendation research at the data level.
@@ 3.5.2. Healthiness Evaluation of Dataset
2) Users prefer less healthy meals in the MealRec dataset. Across the two datasets and two nutritional standards, it is evident that the healthiness of meals favored by users is lower than that of overall meals. For example, in MealRec, the mean FSA score of users, 9.3347, exceeds that of the meals, 9.2406. Existing models usually learn from user historical data to make recommendations. Without the guidance of healthiness in meal recommendations, the model will follow the less healthy meal preferences of some users, which will have a negative impact on the long-term health of these users. This less healthy recommendation phenomenon also emphasizes the necessity of health-aware meal recommendations.
@@ 4. Task Definition
Although MealRec is also suitable for meal plan recommendation, we focus on recommending predefined meals because their research is relatively lagging but they have great application value in reality. The impact of other factors (such as user or meal description, culture, religion, tradition, price, etc.) and meal plan recommendation task are left as future work. The task definition of meal recommendation in this paper can be formulated as follows:
@@ 8. Conclusion and Future Directions
In this paper, we have introduced MealRec, a meal recommendation dataset including meal-course affiliation. MealRec fills the data gap and provides conditions for exploring advanced user preference representation technologies in meal recommendation research. We experiment with different separate and cooperative interaction learning methods and potential ways of improving healthiness, and establish baseline performances in terms of personalization and healthiness. This paper provides new resource and useful insights for AI for good in food domain. In the future, an important direction is to propose user interaction learning methods more suitable for meal recommendation contexts, such as category constraint characteristics of meals (Li et al., 2023a, 2024, b) and other factors like user and meal descriptions (Kawano and Yanai, 2014; Zhou et al., 2024), ingredients (Gao et al., 2022), religion (Sarri et al., 2006), availability (Silva et al., 2018; Wang et al., 2021a), etc. Another direction worthy of research is how to trade off the personalization and healthiness, which is very essential for users to accept healthy meals and develop healthy eating habits. Potential ways are to guide the training of the model in a health-oriented way (Li et al., 2024; Chen et al., 2020), or to treat less healthy recommendations of the model as a bias and improve the health of the recommendations through debias methods (Elsweiler et al., 2017; Chen et al., 2021, 2023; Rokicki et al., 2017).

### v2

# MealRec+: A Meal Recommendation Dataset with Meal-Course Affiliation for Personalization and Healthiness

@@ Abstract.
Meal recommendation, as a typical health-related recommendation task, contains complex relationships between users, courses, and meals. Among them, meal-course affiliation associates user-meal and user-course interactions. However, an extensive literature review demonstrates that there is a lack of publicly available meal recommendation datasets including meal-course affiliation. Meal recommendation research has been constrained in exploring the impact of cooperation between two levels of interaction on personalization and healthiness. To pave the way for meal recommendation research, we introduce a new benchmark dataset called MealRec+. Due to constraints related to user health privacy and meal scenario characteristics, the collection of data that includes both meal-course affiliation and two levels of interactions is impeded. Therefore, a simulation method is adopted to derive meal-course affiliation and user-meal interaction from the user’s dining sessions simulated based on user-course interaction data. Then, two well-known nutritional standards are used to calculate the healthiness scores of meals. Moreover, we experiment with several baseline models, including separate and cooperative interaction learning methods. Our experiment demonstrates that cooperating the two levels of interaction in appropriate ways is beneficial for meal recommendations. Furthermore, in response to the less healthy recommendation phenomenon found in the experiment, we explore methods to enhance the healthiness of meal recommendations. The dataset is available on GitHub (https://github.com/WUT-IDEA/MealRecPlus).
@@ 1. Introduction
To pave the way for meal recommendation studies, we introduce a meal recommendation benchmark dataset called MealRec+. Specifically, considering the category-constraints feature of meals, a mapping from course name phrases to course categories is established. Then, inspired by research on conversational recommender systems, meal-course affiliation and user-meal interaction are derived from user’s dining sessions simulated based on user-course interaction data from a recipe-sharing website222https://www.allrecipes.com. Based on the principle of collaborative filtering that similar users may have similar behaviors, user-meal interactions are mined to expand the scale of data. Finally, two well-known nutritional standards from the World Health Organization and the United Kingdom Food Standards Agency are used to evaluate healthiness of meals. The healthiness analysis of the dataset finds that 1) users have different preferences for healthiness at course-level and meal-level, and 2) users prefer less healthy meals. To explore the impact of different separate and cooperative interaction learning methods on meal recommendation, we experiment with several state-of-the-art models and establish baseline performances. Our experiment demonstrates that cooperating the two levels of interaction in appropriate ways is beneficial for meal recommendations. Furthermore, in response to the less healthy recommendation phenomenon found in the experiment, we explore potential methods to improve the healthiness of recommendation results. In the end, we point out possible directions for building personalized and healthy meal recommender systems, with the long-term goal of transforming how users choose food conveniently and healthily. This aligns with AI for good, viz., utilizing technology’s potential to promote public good health and well-being (Floridi et al., 2021a; Tomašev et al., 2020; Floridi et al., 2021b).
Our main contributions lie two-fold: 1) We introduce a new meal recommendation benchmark dataset, MealRec+, which fills a data gap by including meal-course affiliation and healthiness information. It provides conditions for researching cooperative interaction learning methods on meal recommendation, and serves as a valuable resource for AI for Good research in the food domain. 2) Our experimental study investigates separate and cooperative interaction learning methods and potential ways for improving the healthiness in meal recommendation, which provides useful insights into diet-based public health promotion.
@@ 2.2. Meal Recommendation Datasets
In general, an extensive literature review demonstrates a lack of publicly available meal recommendation datasets with meal-course affiliation. The main reason is that the two main data collecting methods are hindered due to user health privacy and meal scenario characteristics. To create conditions for advancing representation learning in meal recommendation research, we construct a new meal recommendation dataset called MealRec+. A simulation approach is adopted to construct MealRec+, which is inspired by the method of simulating user dialogues to build datasets in conversational recommender system research (Sun and Zhang, 2018; Qu et al., 2023; Tu et al., 2022).
@@ 3. Construction of the Dataset
In order to find the base dataset for constructing our dataset, we search extensively in the literature and the Internet. Finally, we find two open-source large-scale food recommendation datasets, foodRecSys-V1444https://www.kaggle.com/elisaxxyga/foodrecsysv1 and Food180K555https://www.kaggle.com/datasets/shuyangli94/food-com-recipes-and-user-interactions, which come from two very popular food sharing websites2,666https://www.food.com respectively. These two datasets have similar characteristics, such as having similar data sizes and rich course description information. The main difference is that the data density of foodRecSys-V1 is three times that of Food180K, and Food180K contains category tags, but foodRecSys-V1 does not. In our practice, data density is crucial to the construction of meal datasets, so we finally choose foodRecSys-V1 as our base dataset and use the course category tags in Food180K to build a mapping from course name to course category. foodRecSys-V1 contains 49,698 courses and 3,794,003 user-course interactions. The structure of the base dataset is shown in Table 2. The course data contains rich descriptive information including nutrition. The interaction data includes timestamps of each user’s reviews on courses.
@@
timestamp | the time last modified | 2014-04-25T14:54:20 |
@@ 3.1. Course Category Mapping
Based on the mapping table between course name phrases and categories, some courses in the foodRecSys-V1 dataset are assigned categories. Before constructing meal data, we need to clarify the composition of a meal. The basic meal1 comprises three courses: an appetizer, a main dish, and a dessert. In this paper, we refer to the composition of the three-course meal as these three categories have the most courses in the base dataset.
@@ 3.4. Healthiness Measuring
The courses in the base datatset come from the Allrecipes.com website and contain detailed nutritional data powered by the ESHA Research Database 777https://esha.com/. The data format of nutritional data is shown in Figure 3. These nutritional data, quantified per serving, cover all macro-nutrients mentioned in two nutritional standards above. We calculate both the FSA score and WHO score for each course. Following other food-related studies (Trattner and Elsweiler, 2017b), the healthiness of a meal is defined as the mean of FSA/WHO scores of its courses. Thus far, we have constructed the dataset, which called as MealRec+.
@@ 3.5.1. Data Statistics
Compared with other meal recommendation datasets in Table 1, MealRec+ has both levels of interaction and meal-course affiliation, which can support the research of cooperative interaction learning methods. A bundle is a set of items as a whole (Zhu et al., 2014), such as a clothing outfit, a book list, a music playlist, etc. A meal is a type of bundle in the food domain. Research on bundle recommendation (Chang et al., 2020; Ma et al., 2022; Ke et al., 2023a) has emerged in recent years, and Clothing (Sun et al., 2022) and Youshu (Chen et al., 2019) are important benchmark datasets. However, they can’t be applied directly to meal recommendations due to the lack of healthiness information. Table 4 shows the comparison in data size between them. Since Youshu is one of the most important datasets of the same type, we intentionally construct the data size of MealRec+ to be comparable to Youshu. Specifically, the length of the simulated user session affects the size of the constructed dataset, which is chosen to start with a short time and gradually extend until the data size is comparable to Youshu. Furthermore, Clothing and Youshu are used by existing bundle recommendation research to experiment with data-driven deep learning models. Therefore, this comparison of data size shows that MealRec+ can support the applications of existing advanced recommendation technologies to meal recommendation research at the data level.
@@ 3.5.2. Healthiness Evaluation of Dataset
2) Users prefer less healthy meals in the MealRec+ dataset. Across the two datasets and two nutritional standards, it is evident that the healthiness of meals favored by users is lower than that of overall meals. For example, in MealRec, the mean FSA score of users, 9.3347, exceeds that of the meals, 9.2406. Existing models usually learn from user historical data to make recommendations. Without the guidance of healthiness in meal recommendations, the model will follow the less healthy meal preferences of some users, which will have a negative impact on the long-term health of these users. This less healthy recommendation phenomenon also emphasizes the necessity of health-aware meal recommendations.
@@ 4. Task Definition
Although MealRec+ is also suitable for meal plan recommendation, we focus on recommending predefined meals because their research is relatively lagging but they have great application value in reality. The impact of other factors (such as user or meal description, culture, religion, tradition, price, etc.) and meal plan recommendation task are left as future work. The task definition of meal recommendation in this paper can be formulated as follows:
@@ 8. Conclusion and Future Directions
In this paper, we have introduced MealRec+, a meal recommendation dataset including meal-course affiliation. MealRec+ fills the data gap and provides conditions for exploring advanced user preference representation technologies in meal recommendation research. We experiment with different separate and cooperative interaction learning methods and potential ways of improving healthiness, and establish baseline performances in terms of personalization and healthiness. This paper provides new resource and useful insights for AI for good in food domain. In the future, an important direction is to propose user interaction learning methods more suitable for meal recommendation contexts (Zou et al., 2020), such as category constraint characteristics of meals (Li et al., 2023a, 2024, b) and other factors like user and meal descriptions (Kawano and Yanai, 2014; Zhou et al., 2024), ingredients (Gao et al., 2022), religion (Sarri et al., 2006), availability (Silva et al., 2018; Wang et al., 2021a), etc. It is interesting to experiment with other frameworks and recommendation methods (Wang et al., 2020; Huang et al., 2021; Xia et al., 2022). Another direction worthy of research is how to trade off the personalization and healthiness, which is very essential for users to accept healthy meals and develop healthy eating habits. Potential ways are to guide the training of the model in a health-oriented way (Li et al., 2024; Chen et al., 2020), or to treat less healthy recommendations of the model as a bias and improve the health of the recommendations through debias methods (Elsweiler et al., 2017; Chen et al., 2021, 2023; Rokicki et al., 2017).
@@ Acknowledgements.
- Huang et al. (2021) Chao Huang, Jiahui Chen, Lianghao Xia, Yong Xu, Peng Dai, Yanqing Chen, Liefeng Bo, Jiashu Zhao, and Jimmy Xiangji Huang. 2021. Graph-Enhanced Multi-Task Learning of Multi-Level Transition Dynamics for Session-based Recommendation. In Thirty-Fifth AAAI Conference on Artificial Intelligence (AAAI 2021). AAAI Press, 4123–4130.
@@
- Wang et al. (2020) Pengfei Wang, Yu Fan, Long Xia, Wayne Xin Zhao, Shaozhang Niu, and Jimmy X. Huang. 2020. KERL: A Knowledge-Guided Reinforcement Learning Model for Sequential Recommendation. In Proceedings of the 43rd International ACM SIGIR conference on research and development in Information Retrieval (SIGIR ’20). ACM, 209–218.
@@
- Xia et al. (2022) Lianghao Xia, Chao Huang, Yong Xu, Jiashu Zhao, Dawei Yin, and Jimmy X. Huang. 2022. Hypergraph Contrastive Collaborative Filtering. In The 45th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR 2022). ACM, 70–79.
@@
- Zou et al. (2020) Lixin Zou, Long Xia, Yulong Gu, Xiangyu Zhao, Weidong Liu, Jimmy Xiangji Huang, and Dawei Yin. 2020. Neural Interactive Collaborative Filtering. In Proceedings of the 43rd International ACM SIGIR conference on research and development in Information Retrieval (SIGIR ’20). ACM, 749–758.